OpenStudy (anonymous):

evaluate the limit as x goes to 0: ((1-sec^2(6x))/(8x)^2

OpenStudy (anonymous):

I got -9/16

OpenStudy (amistre64):

sin(6x) 1 -1 ------ --------- cos(6x) 8x 8x

OpenStudy (amistre64):

do i see that right?

OpenStudy (anonymous):

uncertain, I think I am right though :-)

OpenStudy (amistre64):

sin(6x) sin(6x) 1 -1 ------ --------- cos(6x) cos(6x) 8x 8x

OpenStudy (amistre64):

the trick is to get 6x under the sin(6x) to make it go to 1

OpenStudy (amistre64):

multiply by 6x.6x/6x.6x and the 8x parts vanish out an x ..

OpenStudy (anonymous):

do you have a link where I can learn l'hopital on my own?

OpenStudy (anonymous):

I just used wikipedia and that confused me :-)

OpenStudy (amistre64):

\[\frac{sin(6x)}{6x}\frac{sin(6x)}{6x}\frac{6}{8}\frac{6}{8}\frac{1}{cos(x)}\frac{-1}{cos(x)}\] right?

OpenStudy (anonymous):

thanks a ton!!!

OpenStudy (amistre64):

lhopitals is just derive the top and the bottom seperately and see what you get when you put in the limit

OpenStudy (anonymous):

gotcha

myininaya (myininaya):

oops missed my negative

OpenStudy (anonymous):

-sin(6x)/6x sin(6x)/6x 6/8cos(6x) 6/8cos(6x)

OpenStudy (anonymous):

yes you did :-)

OpenStudy (amistre64):

and a -1 since (1-sec^2) is equal to -tan^2

OpenStudy (anonymous):

but, you, you;re good!

OpenStudy (amistre64):

i see it there now lol

OpenStudy (anonymous):

yes, that;s what I did